what is the slope-intercept form of the line that contains the following points: (1, –3) and (3, –5).


Answer by checkley71(8403)

(y2-y1)/(x2-x1)=slope thus
(-5+3)/(3-1)
-2/2
-1 is the slope.
now substitute either x & y values snd solve for the y intercept (b) in the line formula
y=mx+b
-5=-1*3+b
-5=-3+b
b=-5+3
b=-2 so the line formula is
y=-1x-2